What mass of solute is present in 40 mL of 6.0mol/L sulfuric acid (H2SO4) solution?
1
Expert's answer
2019-01-10T04:04:58-0500

V(H2SO4)=40ml=0,04L

CM=6 mol/L

m(H2SO4)-?


n=Cm*V= 6*0.04=0.24 mol

M(H2SO4)=98 g/mol

m= M*n=98*0.24=23,52 g



Answer:m=23,52 g